Circular No. 7205

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1999co AND 1999cp J. Y. King and W. D. Li, University of California at Berkeley, on behalf

of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report their discovery of two apparent supernovae on unfiltered CCD images taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 1999co was discovered on frames taken on June 18.43 and 18.47 UT (mag ∼ 17.4) and confirmed on an earlier image taken on June 13.5 (mag ∼ 17.8); an image of the field taken on June 7.5 under poor seeing conditions already showed a hint of the supernova, which is on the northeast shoulder of the faint host galaxy at α = 21^{h}03^{m}19^{s}.12, δ = −14^{o}32′49′′.0 (equinox 2000.0). KAIT images of the same field taken on May 29.5 (limiting mag ∼ 18.0) and 19.6 (limiting mag ∼ 19.0) show nothing at the position of SN 1999co. SN 1999cp was found on CCD frames taken on June 18.2 and 18.3 UT (mag ∼ 18.2), and confirmed on images taken on June 19.2 (mag ∼ 17.2); it is located at α = 14^{h}06^{m}31^{s}.3, δ = −5^{o}26′49′′ (equinox 2000.0), which is ∼ 52′′ west and 23′′ north of the nucleus of NGC 5468. KAIT images of the same field taken on June 13.2 (limiting mag ∼ 19.0) and May 14.3 (limiting mag ∼ 19.5) show nothing at the position of SN 1999cp.

V382 VELORUM K. Mukai, NASA Goddard Space Flight Center and Universities Space

Research Association; and M. Ishida, Institute of Space and Astronautical Sciences, report on behalf of the ASCA team: “ASCA observed V382 Vel between June 9.578 and 10.667 UT (∼ 17 days past optical maximum) for ∼ 40 000 s on-source. We detect the source at ∼ 0.16 counts/s per sensor. The spectrum is relatively hard and highly absorbed, with N_{H} ∼ 9 × 10^{22} cm^{−2} and a bremsstrahlung temperature of ∼ 15 keV. However, no lines are detected (a plasma-model fit requires an abundance of iron ∼ 0.1 that of the sun). A non-thermal origin of (part of) the x-ray flux or a non-ionization equilibrium are both possible explanations for this. The flux in the band 2–10 keV is 2.2 × 10^{−11} erg cm^{−2} s^{−1}; the inferred unabsorbed flux is 3.9 × 10^{−11} erg cm^{−2} s^{−1}, corresponding to apparent and true luminosities of 1.0 × 10^{34} and 1.7 × 10^{34} erg/s, respectively, for an assumed distance of 2 kpc, significantly higher than the BeppoSAX measurements reported by Orio et al. (IAUC 7196). However, we do not observe significant variability within the ∼ 1-day span of the ASCA observation.”

1999 June 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green